# AI-Empowered Integrated Sensing and Communications

### v1

@@ VI-B Case Study II: Hybrid Beamforming Using Learned
We start by presenting the alternating optimization-based method from [148] for designing hybrid beamforming (HBF), which incorporates a modified deep unrolling method. Specifically, it leverages a detection network (DetNet) [149, 150] to unroll the iterations of the projected gradient descent algorithm into a neural network. In addition, we provide a numerical study that evaluates the spectral efficiency performance of the HBF method.
VI-B1 Signal Model
Consider a point-to-point communication system in which the BS and the mobile station are equipped with $N_{t}$ and $N_{r}$ antennas. The post-processed signal at the BS can be expressed as [148]
$\mathbf{y}=\sqrt{\rho}\mathbf{V}^{H}\mathbf{H}\mathbf{F}_{{\mathrm{RF}}}\mathbf{F}_{{\mathrm{BB}}}\mathbf{s}+\mathbf{V}^{H}\mathbf{n},$ (27)
where $\rho$ is the average received power, $\mathbf{V}\in\mathbb{C}^{N_{r}\times N_{s}}$ is the digital combiner, $\mathbf{F}_{\mathrm{RF}}\in\mathbb{C}^{N_{t}\times N_{\mathrm{RF}}}$ is analog beamformer, $\mathbf{F}_{\mathrm{BB}}\in\mathbb{C}^{{N_{\mathrm{RF}}}\times N_{s}}$ is digital baseband precoder, $\mathbf{H}$ is channel matrix, and $\mathbf{n}$ is an AWGN vector with zero mean and variance $\sigma^{2}_{n}$. Here $N_{\mathrm{RF}}$ is the number of RF chains in the BS, $N_{s}$ is the number of data streams, and $\mathbf{s}$ is the signal vector with $\mathbb{E}[\mathbf{s}\mathbf{s}^{H}]=\mathbf{I}$. The achievable spectral efficiency for Gaussian symbols is given by [151]
@@ VI-B1 Signal Model
$\!\!\!\!R=\mathrm{log}_{2}\det\left(\mathbf{I}_{{N_{\mathrm{s}}}}+\frac{\rho}{\sigma_{\mathrm{n}}^{2}N_{\mathrm{s}}}\mathbf{V}^{\dagger}\mathbf{H}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{F}_{\mathrm{BB}}^{H}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}^{H}\mathbf{V}\right),$ (28)
where $(\cdot)^{\dagger}$ is the matrix pseudo-inverse. The spectral efficiency maximization can be approximately achieved by [151]
@@
where $f_{mn}^{\mathrm{RF}}$ is the $(m,n)$th element of $\mathbf{F}_{\mathrm{RF}}$, and $\mathbf{F}_{\mathrm{opt}}$ is the unconstrained optimal digital precoder.
VI-B2 Hybrid Beamforming Design
We now introduce the alternating optimization-based method proposed in [148] for HBF, along with the concept of rolling the projected gradient descent algorithm, inspired by DetNet, which we test here.
The main idea of alternating optimization is to optimize $\mathbf{F}_{\mathrm{RF}}$ with $\mathbf{F}_{\mathrm{BB}}$ given and the second constraint in (29) omitted. Then, $\mathbf{F}_{\mathrm{BB}}$ is optimized to meet the constraint given the optimized $\mathbf{F}_{\mathrm{RF}}$. Thus, the following problem is first considered:

### v2

@@ VI-B Case Study II: Hybrid Beamforming Using Learned
In this case study, we illustrate the algorithm unrolling (also known as algorithm unfolding [15]) approach for designing neural network architectures. To keep the discussion simple and illustrative, we consider an example of hybrid precoding/beamforming design for wireless communications. The extensions to the case of ISAC can be found in [19]. Further extension to robust hybrid precoding/beamforming design for ISAC, where channels are assumed to be inaccurately known and channel mismatches are modeled as additive deterministic norm bounded, can be found in [188].
Figure 9: Unrolled PGA block illustration. The learnable parameters are marked in red color.
VI-B1 System Model
Consider a downlink multi-user MIMO system consisting of an $N$-antenna BS and $K$ single-antenna users. The hybrid beamforming architecture composed of analog beamformer $\mathbf{F}\in\mathbb{C}^{N\times L}$ and digital beamformer $\mathbf{W}=[\mathbf{w}_{1},\ldots,\mathbf{w}_{K}]\in\mathbb{C}^{L\times K}$ is employed for data transmission. Here, $L$ denotes the number of RF chains. The communication data symbol is denoted as $\mathbf{s}=[s_{1},\ldots,s_{K}]\in\mathbb{C}^{K\times 1}$ satisfying $\mathbf{s}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{K})$. The received signal at the $k$-th user side is expressed as
$y_{k}=\mathbf{h}_{k}^{\mathsf{H}}\mathbf{F}\mathbf{w}_{k}s_{k}+\mathbf{h}_{k}^{\mathsf{H}}\sum_{j\neq k}^{K}\mathbf{F}\mathbf{w}_{j}s_{j}+n_{k},$ (29)
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$ denotes the channel between BS and the $k$-th user, and $n_{k}$ is additive white Gaussian noise satisfying $n_{k}\sim\mathcal{CN}(0,\sigma^{2}_{n})$ with $\sigma^{2}_{n}$ being the power of the noise.
From (29), the achievable sum rate is given by
$R=\sum_{k=1}^{K}\mathrm{log}\left(1+\frac{|\mathbf{h}_{k}^{\mathrm{H}}\mathbf{F}\mathbf{w}_{k}|^{2}}{\sum_{j\neq k}^{K}|\mathbf{h}_{k}^{\mathrm{H}}\mathbf{F}\mathbf{w}_{j}|^{2}+\sigma_{n}^{2}}\right).$ (30)
Our goal is to jointly optimize the hybrid beamforming matrices $\mathbf{F}$ and $\mathbf{W}$ to maximize the achievable sum rate $R$. The corresponding optimization problem is formulated as
$\underset{\mathbf{F},\mathbf{W}}{\text{maximize}}\quad$ $R$ (31a)
subject to $|\left[\mathbf{F}\right]_{i,j}|=1,\quad\forall i,j,$ (31b)
$\|\mathbf{F}\mathbf{W}\|_{\text{F}}^{2}=P_{t},$ (31c)
where $|\left[\mathbf{F}\right]_{i,j}|$ denotes the magnitude of the $(i,j)$-th element of matrix $\mathbf{F}$, and $P_{t}$ is the power budget of the BS. The problem (31) is a non-convex problem due to the constant-modulus constraint (31b) and unit-norm constraint (31c). We resort to projected gradient ascent (PGA) and the corresponding unrolled PGA algorithms to solve it, as will be elaborated in the next section.
VI-B2 Projected Gradient Ascent Algorithm
We implement PGA in an alternating-optimization manner. Specifically, keeping $\mathbf{W}$ fixed, $\mathbf{F}$ is updated at the $i$-th iteration via
$\bar{\mathbf{F}}^{(i+1)}$ $=\mathbf{F}^{(i)}+\mu^{(i)}_{\text{F}}\nabla_{\mathbf{F}}R\big|_{\mathbf{F}=\mathbf{F}^{(i)}},$ (32)
$\left[\mathbf{F}^{(i+1)}\right]_{i,j}$ $=\frac{\left[\bar{\mathbf{F}}^{(i+1)}\right]_{i,j}}{\left|\left[\bar{\mathbf{F}}^{(i+1)}\right]_{i,j}\right|},\quad\forall i,j,$ (33)
where $\mu^{(i)}_{\text{F}}$ is the step size for updating $\mathbf{F}$. Similarly, freezing $\mathbf{F}$, $\mathbf{W}$ is updated at the $i$-th iteration via
$\bar{\mathbf{W}}^{(i+1)}$ $=\mathbf{W}^{(i)}+\lambda^{(i)}_{\mathrm{W}}\nabla_{\mathbf{W}}R\big|_{\mathbf{W}=\mathbf{W}^{(i)}},$ (34)
$\mathbf{W}^{(i+1)}$ $=\frac{\sqrt{P_{t}}\bar{\mathbf{W}}^{(i+1)}}{\left\|\mathbf{F}^{(i+1)}\bar{\mathbf{W}}^{(i+1)}\right\|_{\text{F}}},$ (35)
where $\mu^{(i)}_{\text{W}}$ is the step size for updating $\mathbf{W}$. The gradients $\nabla_{\mathbf{F}}R$ and $\nabla_{\mathbf{W}}R$ in (32) and (34) are expressed as [189, 190]
@@ VI-B2 Projected Gradient Ascent Algorithm
$\nabla_{\mathbf{F}}R$ $=\frac{1}{\ln 2}\sum_{k=1}^{K}\left(\frac{\tilde{\mathbf{H}}_{k}\mathbf{FV}}{\text{tr}\left(\mathbf{Z}\tilde{\mathbf{H}}_{k}\right)+\sigma_{n}^{2}}-\frac{\tilde{\mathbf{H}}_{k}\mathbf{F}\mathbf{V}_{\bar{k}}}{\text{tr}\left(\mathbf{Z}_{\bar{k}}\tilde{\mathbf{H}}_{k}\right)+\sigma_{n}^{2}}\right),$ (36)
$\nabla_{\mathbf{W}}R$ $=\frac{1}{\ln 2}\sum_{k=1}^{K}\left(\frac{\bar{\mathbf{H}}_{k}\mathbf{W}}{\text{tr}\left(\mathbf{V}\bar{\mathbf{H}}_{k}\right)+\sigma_{n}^{2}}-\frac{\bar{\mathbf{H}}_{k}\mathbf{W}_{k}}{\text{tr}\left(\mathbf{V}_{\bar{k}}\bar{\mathbf{H}}_{k}\right)+\sigma_{n}^{2}}\right),$ (37)
where $\mathbf{Z}=\mathbf{F}\mathbf{V}\mathbf{F}^{\mathsf{H}}\in\mathbb{C}^{N\times N}$, $\mathbf{Z}_{\bar{k}}=\mathbf{F}\mathbf{V}_{\bar{k}}\mathbf{F}^{\mathsf{H}}\in\mathbb{C}^{N\times N}$, $\mathbf{V}=\mathbf{W}\mathbf{W}^{\mathsf{H}}\in\mathbb{C}^{L\times L}$, $\mathbf{V}_{\bar{k}}=\mathbf{W}_{\bar{k}}\mathbf{W}^{\mathsf{H}}_{\bar{k}}\in\mathbb{C}^{L\times L}$, $\tilde{\mathbf{H}}_{k}=\mathbf{h}_{k}\mathbf{h}^{\mathsf{H}}_{k}\in\mathbb{C}^{N\times N}$, $\bar{\mathbf{H}}_{k}=\mathbf{F}^{\mathsf{H}}\tilde{\mathbf{H}}_{k}\mathbf{F}\in\mathbb{C}^{L\times L}$. Here, $\mathbf{W}_{\bar{k}}$ is obtained by replacing the $k$-th column of $\mathbf{W}$ by an all-zero vector.
VI-B3 Unrolled PGA-based Design
The PGA algorithm is capable of optimizing the hybrid transmit beamforming matrices $\mathbf{F}$ and $\mathbf{W}$ within a predefined number of layers. However, the key challenge in PGA algorithm stems from determining the optimal hyperparameters (i.e., step sizes $\mu_{\text{F}}$ and $\mu_{\text{W}}$ in our context), which can significantly affect the performance and convergence of the algorithm [191]. To this end, we resort to algorithm unrolling technique [192] to tune these hyperparameters in an unsupervised learning manner. The main principle of this technique is to transform each iteration of PGA algorithm into a network layer and stack these layers together [192, 191], resulting a multi-layer neural network, as shown in Fig. 9. In the unrolled network, the trainable parameters, i.e., the step sizes for $I$ iterations, are defined as $\bm{\Phi}=[\bm{\mu}_{1},\ldots,\bm{\mu}_{I}]$, where $\bm{\mu}_{i}=[\mu^{(i)}_{\text{F}},\,\mu^{(i)}_{\text{W}}]^{\mathsf{T}}$ stands for the step size vector to be learned for the $i$-th layer. To train the unrolled network, we employ a loss function that takes into account the outputs across all intermediate layers, expressed as [193]
$\mathcal{L}(\bm{\Phi})=-\frac{1}{|\mathcal{D}|}\sum^{|\mathcal{D}|}_{d=1}\frac{1}{I}\sum^{I}_{i=1}\mathrm{log}\left(1+i\right)R(\mathbf{h}^{(d)}_{k},\mathbf{F}^{(i)},\mathbf{W}^{(i)})$ (38)
in which $\mathcal{D}$ is the data set containing communication channel realizations. Then the hyperparameter matrix $\bm{\Phi}$ is tuned via
$\bm{\Phi}^{\star}=\mathrm{arg}\min_{\bm{\Phi}}\mathcal{L}(\bm{\Phi}).$ (39)
VI-B4 Simulation
In this section, the performance of unrolled neural network tailored to hybrid beamforming design is tested. The system parameters are set as $N=16$, $L=6$, and $K=4$. The noise variance is set as $\sigma^{2}_{n}=1$. The SNR is defined as $\text{SNR}=10\mathrm{log}_{10}(P_{t}/\sigma^{2}_{n})$. The channel $\{\mathbf{h}_{k}\}^{K}_{k=1}$ is modeled as Rayleigh fading channel. The sizes of data sets for training and testing are 1000 and 100. For training phase, stochastic gradient descent optimizer with a learning rate of 0.005 is used. The fixed step sizes for the PGA and initial step sizes for unrolled PGA are both set as 0.05. In Figs. 10 and 11, the hybrid beamforming design using unrolled neural network is named as ‘unrolled PGA’, and PGA-based hybrid beamforming design as ‘PGA’.
The achievable sum rate versus the number of iterations/layers for $\text{SNR}=10\,\text{dB}$ is shown in Fig. 10. Both the PGA and unrolled PGA are observed to stabilize at distinct constant values, demonstrating the strong convergence of the proposed schemes. Notably, the unrolled PGA converges much faster than the benchmark PGA, highlighting the effectiveness of algorithm unrolling in optimizing the hybrid transmit beamformer. The reason for unrolled PGA to outperform the PGA is that the unrolled PGA algorithm is capable of learning the update rules from data, allowing it to fit the objective function. This capacity allows the unrolled PGA algorithm to escape local optima more effectively than the PGA [188].
Figure 10: Achievable sum rate versus number of iterations/layers with $\text{SNR}=10~\text{dB}$, $N=16$, $K=4$, and $L=6$.
Figure 11: Achievable sum rate versus SNR with $N=16$, $L=6$, and $K=4$.
Figure 11 shows the achievable sum rate versus SNR varying from 5-5 to $10$ dB. It is observed that the unrolled PGA consistently achieves a higher sum rate than the standard PGA across the entire SNR range. This implies that the unrolled PGA maintains superior performance over the benchmark PGA even at low SNR levels.
VI-C Case Study III: Autoencoder-Based ISAC Constellation Design
A key enabler in realizing the promise of ISAC is the development of waveforms capable of supporting both information transmission and target sensing. In this context, the design of modulation constellations plays a pivotal role, as it directly influences the performance of both communication and sensing functionalities.
@@ VI-C Case Study III: Autoencoder-Based ISAC Constellation Design
Figure 12: Block diagram of an ISAC system consisting of an ISAC base station (transmitter and radar), a UE communication receiver, and a car as the radar target. The objective is to transmit a message $\mathbf{m}$ to UE and detect the presence of the target using a single transmitted signal $x$. The received signals after passing through the communication and radar channels are denoted by $y$ and $z$, respectively, where $y\sim f_{Y|X}(y|x)$ and $z\sim f_{Z|X}(z|x)$. The neural networks $\Psi_{E}$, $\Psi_{\rm comm}$, and $\Psi_{\rm radar}$ represent the encoder, communication decoder, and radar detector.
A fundamental challenge in ISAC arises from the inherent trade-off between the randomness that benefits communication and the deterministic structure preferred for accurate sensing. Traditional constellations typically offer either high communication efficiency or high sensing accuracy—but not both. Specifically, QAM is optimized for communication, while PSK is more suitable for sensing [194, 70, 195]. Thus, constellation shaping has emerged crucial for balancing sensing and communication requirements in ISAC systems, and forming a unified waveform.
Waveform design in ISAC is typically formulated as an optimization problem that balances communication and sensing objectives, such as maximizing data rate and minimizing BER for communication, while improving range and velocity resolution for sensing, through a joint cost function. Specifically, optimization problem (2) can incorporate meaningful sensing and communication metrics for waveform design. In general, the metrics listed in Table III and Table IV may be used. It is often desirable to design a waveform that simultaneously minimizes the BER for communication and maximizes $P_{D}$ for radar.
In this case study, we present an example of how such an optimization problem can be solved using machine learning. Before presenting the design details, we first describe the problem and the system model.
Consider the ISAC system illustrated in Fig. 12, where, at each channel use, the ISAC base station aims to transmit a binary message vector $\bm{m}\in\{0,1\}^{K}$ and simultaneously sense a target using a single transmitted signal. Since each input $\bm{m}$ is a binary vector of length $K$, there are $M=2^{K}$ possible distinct message vectors. Traditional approaches may use an $M$-ary constellation, such as $M$-PSK, to map each message to a constellation point $x(\bm{m})\in\mathbb{C}$ and transmit it.
The transmitted signal $x(\bm{m})\in\mathbb{C}$ passes through both the communication channel $f_{Y|X}(y|x)$ and the sensing channel $f_{Z|X}(z|x)$. Assuming that both channels are AWGN channels, the communication and sensing signals are respectively given by
$y$ $=x(\bm{m})+n_{\rm comm},$ (40a)
$z$ $=x(\bm{m})+n_{\rm radar},$ (40b)
where $n_{\rm comm}\sim\mathcal{CN}(0,\sigma^{2}_{c})$ and $n_{\rm radar}\sim\mathcal{CN}(0,\sigma^{2}_{\rm radar})$ are the complex Gaussian noises associated with each channel, respectively.
ISAC research [194, 70, 195], shows that constellation structure plays a key role in balancing performance trade-offs between these two functions. While QAM-like constellations favor communication and constant modulus constellations like PSK benefits sensing. Hence, intermediate designs like multi-ring constellations offer a promising balance, and thus are more suited for joint sensing and communication.
In the following, we illustrate how a combination of supervised and unsupervised learning can be used to design effective constellations for ISAC.
VI-C1 Communication
The goal of communication is to send a binary message vector $\bm{m}$ such that it can be reconstructed at the UE with minimal errors. To enable this, we can design an end-to-end communication system based on an autoencoder architecture. In Fig. 12, the encoder and decoder neural networks are denoted by $\Psi_{\rm E}$ and $\Psi_{\rm comm}$, respectively.
To minimize the BER, the autoencoder must learn to generate $M=2^{K}$ well-separated clusters in the latent space. Thus, the autoencoder effectively designs an $M$-ary constellation. For an AWGN channel, a well-trained autoencoder should maximize the minimum distance between constellation points to improve robustness against noise. At each channel use, one of these constellation points is then transmitted.
The autoencoder-based end-to-end design framework has been successfully applied to various communication scenarios (e.g., [30, 156, 196]). The objective is to train a DNN that accurately estimates the transmitted bits, i.e., to minimize the number of errors. To this end, the cross-entropy loss function is commonly used to guide the training process.
Let $\bm{m}_{j}$ denote the $j$th message bit, and let $\hat{\bm{m}}_{j}$ represent the corresponding likelihood estimate produced by the neural network $\Psi_{\rm comm}$. The cross-entropy loss is then defined as
$\mathcal{L}_{\rm comm}=-\mathbb{E}\Big[\sum_{j=1}^{K}\bm{m}_{j}\mathrm{log}\left(\hat{\bm{m}}_{j}\right)\Big],$ (41)
where $\mathbb{E}$ denotes the expectation operator.
VI-C2 Sensing
For sensing, we assume a simple task where the BS acts as a radar and wants to detect a target, as shown in Fig. 12. Assume that that the target be present ($T=1$) or absent ($T=0$) with equal probability, i.e., $p(T=1)=0.5$.
For radar, we use a supervised approach to learn the model $\Psi_{\rm radar}$ by defining a loss function for target detection. One may also define loss functions for other metrics, such as angle estimation. Our loss function for target detection is the binary cross-entropy loss function, given by